OpenStudy (anonymous):

x^2 - 6x + 8=(x +4 )( x - 2) right?

OpenStudy (apoorvk):

niet... simplify the RHS to see if you get LHS..

OpenStudy (anonymous):

huhh?

OpenStudy (anonymous):

\[x^2+2 x-8=(x+4)(x-2) \]

OpenStudy (anonymous):

isnt that what i put?

OpenStudy (campbell_st):

its (x-4)(x-2) = x^2 -6x + 8